American products plans to lay off 10% of its employees in its aerospace divison and 15 % of its employees in its argicultural division. If altogether 12 % of the 3000 employees in these two divison will be laid off, then how many employees are in each division?
Answer by checkley75(3666)

.10X+.15(3000-X)=.12*3000
.10X+450-.15X=360
-.05X=360-450
-.05X=-90
X=-90/-.05
X=1800 IS THE AMOUNT OF EMPLOYEES IN AEROSPACE.
3000-1800=1200 AMOUNT OF EMPLOYEES IN AGRICULTURE.
PROOF
.10*1800+.15*1200=360
180+180=360
360=360